OpenStudy (anonymous):

Explain how the Quotient of Powers was used to simplify this expression. 5^4/25 = 5^2 @mathmale

OpenStudy (anonymous):

Would you like the question's options?

OpenStudy (mathmale):

I did hope you'd use my suggestion (to use parentheses for additional clarity). Please re-write your expression so I know whether you mean (5^4) / 25 or 5^(4/25).

OpenStudy (anonymous):

Sorry it's (5^4)/25 = (5^2)

OpenStudy (anonymous):

I think you're supposed to make 25 into an exponent and then subtract the exponents to get (5^2)

OpenStudy (mathmale):

Huge difference in clarity. Thank you. \[\frac{ 5^4 }{ 25 }=?\] Do you see any way of reducing this by manipulating the base and exponent?

OpenStudy (mathmale):

Leave that 5^4 as is. Rewrite 25 as a power of 5. Then follow your own suggestion regarding the exponents.

OpenStudy (anonymous):

So you're supposed to subtract the exponents in the end?

OpenStudy (mathmale):

Exactly!

OpenStudy (anonymous):

Oh ok thank you so much !

OpenStudy (mathmale):

\[\frac{ x^a}{ x^b }=x ^{a-b}\]

OpenStudy (mathmale):

I ask for a bit more effort up front, so that you eliminate ambiguity, but trust me, it's worth being crystal clear when you're communicating math. Happy to work with you.

OpenStudy (anonymous):

Same thanks again